Problem 53 (modified): 60 students in your class are going to be split into 5 scribing teams of 12. Unknown to all of you is there is A whom both B and C fancy. If we pick the teams at random (instead of alphabetically), how many different ways will (i) both B and C land up with A in the same team, (ii) neither B or C land up in a team with A, (iii) exactly one of B and C lands up with A on a team?
Answer #1

Solution :-

Given data:

Number of students = 60

There are 5 teams, each team consists of 12 students.

Consider the 5 teams as 1, 2, 3, 4, 5. and A, B, and C are three members.

( i ) :-

Here we need to find out the number of different ways will both B and C land up with A in the same team.

There are 5 teams, so in 5 different ways both B and C land up with A in the same team.

\therefore In 5 different ways both B and C land up with A in the same team.

( ii ) :-

Here we need to find out the number of different ways will neither B or C land up in a team with A.

Now consider, If A is in team-1, then B and C are either in team-2, team-3, team-4 and team-5. If A is in team-2, then B and C are either in team-1, team-3, team-4 and team-5. and so on.

i.e.

if A is in team-1 then B,C is in one of these teams 2, 3, 4, 5.

A is in team-2 then B,C is in one of these teams 1, 3, 4, 5.

A is in team-3 then B,C is in one of these teams 1, 2, 4, 5.

A is in team-4 then B,C is in one of these teams 1, 2, 3, 5.

A is in team-5 then B,C is in one of these teams 1,2, 3, 4.

Hence, The total number of ways = 4 * 5

= 20

\therefore The number of different ways will neither B or C land up in a team with A is 20 ways.

( iii ) :-

Here we need to find out the number of different ways will exactly one of B and C lands up with A on a team.

Here consider the number of possibilities. they are

if A & C is in team-1 then B is in one of these teams 2, 3, 4, 5.

A & C is in team-2 then B is in one of these teams 1, 3, 4, 5.

A & C is in team-3 then B is in one of these teams 1, 2, 4, 5.

A & C is in team-4 then B is in one of these teams 1, 2, 3, 5.

A & C is in team-5 then B is in one of these teams 1,2, 3, 4.

Hence, number of possibilities are 4 * 5 = 20 ways.

if A & B is in team-1 then C is in one of these teams 2, 3, 4, 5.

A & B is in team-2 then C is in one of these teams 1, 3, 4, 5.

A & B is in team-3 then C is in one of these teams 1, 2, 4, 5.

A & B is in team-4 then C is in one of these teams 1, 2, 3, 5.

A & B is in team-5 then C is in one of these teams 1,2, 3, 4.

Hence, number of possibilities are 4 * 5 = 20 ways.

Hence the Total number of ways = 20 + 20

= 40 ways.

\therefore The number of different ways will exactly one of B and C lands up with A on a team​​​​​​​ is 40 ways.